# Convergence of a Finite Volume Scheme for Compactly Heterogeneous Scalar Conservation Laws

### v2

@@ 3.3.1 Compensated compactness
$|Q_{2,1}|$ $=|H(x_{j},\rho_{j+1/2}^{n,-})-H(x_{j},\rho_{j+1/2}^{n,+})|$
$\leq\underbrace{|H(x_{j},\rho_{j+1/2}^{n,-})-H(x_{j+1},\rho_{j+1/2}^{n,+})|}_{=0}+|H(x_{j+1},\rho_{j+1/2}^{n,+})-H(x_{j},\rho_{j+1/2}^{n,+})|$
$\leq\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{x}H(x,p)|\;\Delta x.$
Otherwise, assume to fix the ideas that $\rho^{-}\leq\alpha_{j}\leq\rho^{+}$. Then, setting $L_{H}\coloneq\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{uu}H(x,p)|$ we write
$Q_{2,1}$ $=\int_{\rho^{-}}^{\rho^{+}}|\partial_{u}H(x_{j},\xi)-\partial_{u}H(x_{j},\alpha_{j})|\mathinner{\mathrm{d}{\xi}}$
$\leq L_{H}\int_{\rho^{-}}^{\rho^{+}}|\xi-\alpha_{j}|\mathinner{\mathrm{d}{\xi}}$
@@
$\sum_{n=0}^{N}\Delta t\sum_{|j|\leq J}\left|[\![Q_{2}(w_{\Delta},u_{\Delta})]\!]_{x_{j+1/2}}^{n}\right|$
$\leq L\sum_{n=0}^{N}\Delta t\sum_{|j|\leq J}\left\{L_{H}\;(\rho_{j+1/2}^{n,-}-\rho_{j+1/2}^{n,+})^{2}+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{x}H(x,p)|\;\Delta x+2(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{xu}^{2}H(x,p)|\Delta x\right\}$
$\leq L\left\{L_{H}+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{x}H(x,p)|+2(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{xu}^{2}H(x,p)|\right\}(\lambda c_{2}+2TR),$
which ensures that $I_{4,2}(\phi)$ belongs to a compact subset of $\mathbf{W}^{-1,q_{2}}(\Omega)$.
@@
Assume that $H$ satisfies (CH)–(CVX). Fix $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$ and let $(u_{\Delta})_{\Delta}$ be the sequence generated by the scheme described in Section 2.
Then, there exists a limit function $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ such that along a subsequence as $\Delta\to 0$, $(u_{\Delta})_{\Delta}$ converges in $\mathbf{L}_{\mathbf{loc}}^{p}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ for all $p\in[1,+\infty\mathclose{[}$ and a.e. on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$ to $u$.
@@ 3.3.2 Convergence
To prove that the limit function $u$ of Corollary 3.16 is the entropy solution, we will need the following technical result.
@@
Let us introduce the piecewise constant function: ${\rho_{\Delta}\coloneq\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}u_{j}^{n}\mathbf{1}_{[t^{n},t^{n+1}\mathclose{[}\times I_{j}}}$.
[…]
$|u_{\Delta}(t,x)-\rho_{\Delta}(t,x)|=|\mathcal{U}^{n}(t,x)-u_{j+1}^{n}|\leq|\rho_{j+1/2}^{n,+}-u_{j+1}^{n}|,$
@@
$|u_{\Delta}(t,x)-\rho_{\Delta}(t,x)|=|\mathcal{U}^{n}(t,x)-u_{j}^{n}|\leq|\rho_{j+1/2}^{n,-}-u_{j}^{n}|.$
@@
$\sum_{n=0}^{N}\sum_{|j|\leq K}\int_{t^{n}}^{t^{n+1}}\int_{x_{j}}^{x_{j+1}}|u_{\Delta}(t,x)-\rho_{\Delta}(t,x)|^{2}\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
$\leq\frac{1}{2}\sum_{n=0}^{N}\sum_{|j|\leq K}\left\{|\rho_{j+1/2}^{n,+}-u_{j+1}^{n}|^{2}+|\rho_{j+1/2}^{n,-}-u_{j}^{n}|^{2}\right\}\Delta x\Delta t\leq\frac{c_{1}}{2}\Delta t,$
@@
$\left(|u_{j}^{n+1}-k|-|u_{j}^{n}-k|\right)\Delta x+(\Phi_{j+1/2}^{n}-\Phi_{j-1/2}^{n})\Delta t$ (3.19)
@@
$|u_{j}^{n+1}-H_{j}(k,k,k)|$ $\geq|u_{j}^{n+1}-k+\lambda(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j+1/2}(k,k))|$
$=|u_{j}^{n+1}-k|+\lambda\mathop{\rm sgn}(u_{j}^{n+1}-k)(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k)).$
@@
$|u_{j}^{n+1}-H_{j}(k,k,k)|$ $\leq H_{j}(u_{j-1}^{n}\vee k,u_{j}^{n}\vee k,u_{j+1}^{n}\vee k)-H_{j}(u_{j-1}^{n}\wedge k,u_{j}^{n}\wedge k,u_{j+1}^{n}\wedge k)$
$=|u_{j}^{n}-k|-\lambda(\Phi_{j+1/2}^{n}-\Phi_{j-1/2}^{n}).$
@@
Let $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}^{+})$ fix $T>0$, $R$ such that the support of $\varphi$ is included in $[0,T]\times[-R,R]$. Define
$\forall n\in\mathbb{N},\;\forall j\in\mathbb{Z},\quad\varphi_{j}^{n}\coloneq\frac{1}{\Delta x}\int_{I_{j}}\varphi(t^{n},x)\mathinner{\mathrm{d}{x}}.$
@@
$A$ :-$\coloneq\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}|u_{j}^{n}-k|(\varphi_{j}^{n}-\varphi_{j}^{n-1})\Delta x+\sum_{j\in\mathbb{Z}}|u_{j}^{o}-k|\varphi_{j}^{o}\Delta x:$
$B$ :-$\coloneq\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\Phi_{j+1/2}^{n}(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t:$
$C$ :-$\coloneq-\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\mathop{\rm sgn}(u_{j}^{n+1}-k)(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k))\varphi_{j}^{n}\Delta t.:$
@@
$A{\ \underset{\Delta\to 0}{\longrightarrow}\ }\int_{0}^{+\infty}\int_{\mathbb{R}}|u-k|\partial_{t}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+\int_{\mathbb{R}}u_{o}(x)\varphi(0,x)\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}.$
@@
$\left|\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}(B_{1}+B_{5})(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t\right|$ $\leq 2L\biggl{(}T\|\partial_{xx}^{2}\varphi\|_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1})}+\|\partial_{tx}^{2}\varphi\|_{\mathbf{L}^{1}}\biggr{)}(\Delta x+\Delta t)$
$+2L\int_{0}^{+\infty}\int_{\mathbb{R}}|\rho_{\Delta}(t,x+\Delta x)-\rho_{\Delta}(t,x)|\cdot|\partial_{x}\varphi(t,x)|\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}.$
Since $(\rho_{\Delta})_{\Delta}$ converges a.e. on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$, $(\rho_{\Delta}\partial_{x}\varphi)_{\Delta}$ is strongly compact in $\mathbf{L}^{1}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$. As a consequence of the Riesz-Fréchet-Kolmogorov compactness characterization, see [6, Chapter 4, Section 5],
@@
$\left|\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}(B_{2}+B_{4})(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t\right|\leq 2\mu T\cdot\|\partial_{x}\varphi\|_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1})}\Delta x.$
@@
$\left|\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}B_{3}(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t\right|\leq T\;\sup_{\underset{p,k\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}|\partial_{x}\Phi(x,p,k)|\cdot\|\partial_{x}\varphi\|_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1})}\Delta x.$
@@
$\leq T\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\Phi(x,p,k)|\cdot\|\partial_{tx}^{2}\varphi\|_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1})}\Delta x+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\Phi(x,p,k)|\cdot\|\partial_{tx}^{2}\varphi\|_{\mathbf{L}^{1}}\Delta t$
$+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{x}\Phi(x,p,k)|\cdot\|\partial_{x}\varphi\|_{\mathbf{L}^{1}}\Delta x+\int_{0}^{+\infty}\int_{\mathbb{R}}\Phi(x,\rho_{\Delta},k)\partial_{x}\varphi(t,x)\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}},$
@@
Finally, introduce $F_{\Delta}$ the continuous, linear affine function such that $F_{\Delta}(x_{j+1/2})=F_{\mathrm{int}}^{j+1/2}(k,k)$ for all $j\in\mathbb{Z}$. As a consequence of Lemma 3.7, $(F_{\Delta})_{\Delta}$ converges uniformly to $x\mapsto H(x,k)$ on $\mathbb{R}$. Moreover, since $(F_{\Delta})_{\Delta}$ is bounded in $\mathbf{Lip}(\mathbb{R},\mathbb{R})$, $(F_{\Delta}^{\prime})_{\Delta}$ converges weakly* in $\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$, and the limit can only be $x\mapsto\partial_{x}H(x,k)$. We can now conclude with $C$. Write $C=C_{1}+C_{2}$, where
$C_{1}$ :-$\coloneq-\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}\mathop{\rm sgn}(u_{j}^{n}-k)(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k))(\varphi_{j}^{n-1}-\varphi_{j}^{n})\Delta t:$
$C_{2}$ :-$\coloneq-\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}\int_{I_{j}}\mathop{\rm sgn}(u_{j}^{n}-k)F_{\Delta}^{\prime}(x)\varphi_{j}^{n}\Delta t.:$
@@
$|C_{1}|\leq\left(2\mu+\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\right)\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}\left|\varphi_{j}^{n-1}-\varphi_{j}^{n}\right|\Delta x\Delta t\leq 2TR\left(2\mu+\sup_{x\in\mathbb{R}}|\partial_{x}H(x,k)|\right)\|\partial_{t}\varphi\|_{\mathbf{L}^{\infty}}\Delta t.$
@@
for any $k\in[\underline{u},\overline{u}]\backslash\mathcal{L}$. Then, repeat the argument from [25] to extend it for all $k\in[\underline{u},\overline{u}]$.

### v3

@@ 3.3.1 Compensated compactness
${\left|Q_{2,1}\right|}$ $={\left|H(x_{j},\rho_{j+1/2}^{n,-})-H(x_{j},\rho_{j+1/2}^{n,+})\right|}$
$\leq\underbrace{{\left|H(x_{j},\rho_{j+1/2}^{n,-})-H(x_{j+1},\rho_{j+1/2}^{n,+})\right|}}_{=0}+{\left|H(x_{j+1},\rho_{j+1/2}^{n,+})-H(x_{j},\rho_{j+1/2}^{n,+})\right|}$
$\leq\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{x}H(x,p)\right|}\;\Delta x.$
Otherwise, assume to fix the ideas that $\rho^{-}\leq\alpha_{j}\leq\rho^{+}$. Then, setting $L_{H}\coloneq\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{uu}^{2}H(x,p)\right|}$ we write
$Q_{2,1}$ $=\int_{\rho^{-}}^{\rho^{+}}{\left|\partial_{u}H(x_{j},\xi)-\partial_{u}H(x_{j},\alpha_{j})\right|}\mathinner{\mathrm{d}{\xi}}$
$\leq L_{H}\int_{\rho^{-}}^{\rho^{+}}{\left|\xi-\alpha_{j}\right|}\mathinner{\mathrm{d}{\xi}}$
@@
$\sum_{n=0}^{N}\Delta t\sum_{{\left|j\right|}\leq J}{\left|[\![Q_{2}(\xi_{\Delta},u_{\Delta})]\!]_{x_{j+1/2}}^{n}\right|}$
$\leq L\sum_{n=0}^{N}\Delta t\sum_{{\left|j\right|}\leq J}\left\{L_{H}\;(\rho_{j+1/2}^{n,-}-\rho_{j+1/2}^{n,+})^{2}+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{x}H(x,p)\right|}\;\Delta x+2(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{xu}^{2}H(x,p)\right|}\Delta x\right\}$
$\leq L\left\{L_{H}+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{x}H(x,p)\right|}+2(\overline{u}-\underline{u})\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{xu}^{2}H(x,p)\right|}\right\}(\lambda c_{2}+2TR),$
which ensures that $I_{4,2}(\phi)$ belongs to a compact subset of $\mathbf{W}^{-1,q_{2}}(\Omega,\mathbb{R})$.
@@
Assume that $H$ satisfies (C3)–(CH)–(CVX). Fix $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$ and let $(u_{\Delta})_{\Delta}$ be the sequence generated by the scheme described in Section 2. Then, there exists a limit function $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ such that along a subsequence, as $\Delta\to 0$ while satisfying (2.3), $(u_{\Delta})_{\Delta}$ converges in $\mathbf{L}_{\mathbf{loc}}^{p}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ for all $p\in[1,+\infty\mathclose{[}$ and a.e. on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$ to $u$.
@@ 3.3.2 Convergence
Equipped with compactness obtained in Corollary 3.17, we prove that the limit function is the entropy solution. We will need the following technical result.
@@
Introduce the piecewise constant function: ${\rho_{\Delta}\coloneq\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}u_{j}^{n}\mathbf{1}_{[t^{n},t^{n+1}\mathclose{[}\times I_{j}}}$.
We claim that ${\left\|u_{\Delta}-\rho_{\Delta}\right\|}_{\mathbf{L}_{\mathbf{loc}}^{2}(\mathbb{R}^{+}\times\mathbb{R})}{\ \underset{\Delta\to 0}{\longrightarrow}\ }0$. Indeed, for all $(t,x)\in[t^{n},t^{n+1}\mathclose{[}\times\mathopen{]}x_{j+1/2},x_{j+1}\mathclose{[}$,
${\left|u_{\Delta}(t,x)-\rho_{\Delta}(t,x)\right|}={\left|\mathcal{U}^{n}(t,x)-u_{j+1}^{n}\right|}\leq{\left|\rho_{j+1/2}^{n,+}-u_{j+1}^{n}\right|},$
@@
${\left|u_{\Delta}(t,x)-\rho_{\Delta}(t,x)\right|}={\left|\mathcal{U}^{n}(t,x)-u_{j}^{n}\right|}\leq{\left|\rho_{j+1/2}^{n,-}-u_{j}^{n}\right|}.$
@@
$\sum_{n=0}^{N}\sum_{{\left|j\right|}\leq K}\int_{t^{n}}^{t^{n+1}}\int_{x_{j}}^{x_{j+1}}{\left|u_{\Delta}(t,x)-\rho_{\Delta}(t,x)\right|}^{2}\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
$\leq\frac{1}{2}\sum_{n=0}^{N}\sum_{{\left|j\right|}\leq K}\left\{{\left|\rho_{j+1/2}^{n,+}-u_{j+1}^{n}\right|}^{2}+{\left|\rho_{j+1/2}^{n,-}-u_{j}^{n}\right|}^{2}\right\}\Delta x\Delta t\leq\frac{c_{1}}{2}\Delta t,$
@@
$\left({\left|u_{j}^{n+1}-k\right|}-{\left|u_{j}^{n}-k\right|}\right)\Delta x+(\Phi_{j+1/2}^{n}-\Phi_{j-1/2}^{n})\Delta t$ (3.20)
@@
${\left|u_{j}^{n+1}-H_{j}(k,k,k)\right|}$ $={\left|u_{j}^{n+1}-k+\lambda(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k))\right|}$
$\geq{\left|u_{j}^{n+1}-k\right|}+\lambda\mathop{\rm sgn}(u_{j}^{n+1}-k)(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k)).$
@@
${\left|u_{j}^{n+1}-H_{j}(k,k,k)\right|}$ $\leq H_{j}(u_{j-1}^{n}\vee k,u_{j}^{n}\vee k,u_{j+1}^{n}\vee k)-H_{j}(u_{j-1}^{n}\wedge k,u_{j}^{n}\wedge k,u_{j+1}^{n}\wedge k)$
$={\left|u_{j}^{n}-k\right|}-\lambda(\Phi_{j+1/2}^{n}-\Phi_{j-1/2}^{n}).$
@@
Let $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}^{+})$ and fix $T>0$, $R$ such that the support of $\varphi$ is included in $0,T]\times[-R,R]$. Define for all $n\in\mathbb{N}$ and $j\in\mathbb{Z}$, $\varphi_{j}^{n}\coloneq\frac{1}{\Delta x}\int_{I_{j}}\varphi(t^{n},x)\mathinner{\mathrm{d}{x}}$
@@
$A$ $\coloneq\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}{\left|u_{j}^{n}-k\right|}(\varphi_{j}^{n}-\varphi_{j}^{n-1})\Delta x+\sum_{j\in\mathbb{Z}}{\left|u_{j}^{o}-k\right|}\varphi_{j}^{o}\Delta x$
$B$ $\coloneq\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\Phi_{j+1/2}^{n}(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t$
$C$ $\coloneq-\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}\mathop{\rm sgn}(u_{j}^{n+1}-k)(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k))\varphi_{j}^{n}\Delta t.$
@@
$A{\ \underset{\Delta\to 0}{\longrightarrow}\ }\int_{0}^{+\infty}\int_{\mathbb{R}}{\left|u-k\right|}\partial_{t}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+\int_{\mathbb{R}}{\left|u_{o}(x)-k\right|}\varphi(0,x)\mathinner{\mathrm{d}{x}}.$
@@
${\left|\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}(B_{1}+B_{5})(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t\right|}$ $\leq 2L\bigl(T{\left\|\partial_{xx}^{2}\varphi\right\|}_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1}(\mathbb{R}))}+{\left\|\partial_{tx}^{2}\varphi\right\|}_{\mathbf{L}^{1}(\mathbb{R}^{+}\times\mathbb{R})}\bigr)(\Delta x+\Delta t)$
$+2L\int_{0}^{+\infty}\int_{\mathbb{R}}{\left|\rho_{\Delta}(t,x+\Delta x)-\rho_{\Delta}(t,x)\right|}\cdot{\left|\partial_{x}\varphi(t,x)\right|}\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}.$
Since $(\rho_{\Delta})_{\Delta}$ converges a.e. on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}$, $(\rho_{\Delta}\partial_{x}\varphi)_{\Delta}$ is strongly compact in $\mathbf{L}^{1}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$. As a consequence of the Riesz-Fréchet-Kolmogorov compactness characterization, see [BrezisBook, Chapter 4, Section 5],
@@
${\left|\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}(B_{2}+B_{4})(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t\right|}\leq 2\mu T\cdot{\left\|\partial_{x}\varphi\right\|}_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1}(\mathbb{R}))}\Delta x.$
@@
${\left|\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}B_{3}(\varphi_{j+1}^{n}-\varphi_{j}^{n})\Delta t\right|}\leq T\;\sup_{\underset{p,k\in[\underline{u},\overline{u}]}{x\in\mathbb{R}}}{\left|\partial_{x}\Phi(x,p,k)\right|}\cdot{\left\|\partial_{x}\varphi\right\|}_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1}(\mathbb{R}))}\Delta x.$
@@
$\leq T\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\Phi(x,p,k)\right|}\cdot{\left\|\partial_{tx}^{2}\varphi\right\|}_{\mathbf{L}^{\infty}(\mathbb{R}^{+},\mathbf{L}^{1}(\mathbb{R}))}\Delta x+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\Phi(x,p,k)\right|}\cdot{\left\|\partial_{tx}^{2}\varphi\right\|}_{\mathbf{L}^{1}(\mathbb{R}^{+}\times\mathbb{R})}\Delta t$
$+\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{x}\Phi(x,p,k)|\cdot{\left\|\partial_{x}\varphi\right\|}_{\mathbf{L}^{1}(\mathbb{R}^{+}\times\mathbb{R})}\Delta x+\int_{0}^{+\infty}\int_{\mathbb{R}}\Phi(x,\rho_{\Delta},k)\partial_{x}\varphi(t,x)\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}},$
@@
Finally, introduce $F_{\Delta}$ the continuous, piecewise linear affine function such that $F_{\Delta}(x_{j+1/2})=F_{\mathrm{int}}^{j+1/2}(k,k)$ for all $j\in\mathbb{Z}$. Let us recall that because of (CH), $F_{\Delta}$ is constant outside a compact subset.
As a consequence of Lemma 3.7, $(F_{\Delta})_{\Delta}$ converges uniformly to $x\mapsto H(x,k)$ on $\mathbb{R}$. Moreover, since $(F_{\Delta})_{\Delta}$ is bounded in $\mathbf{Lip}(\mathbb{R},\mathbb{R})$, $(F_{\Delta}^{\prime})_{\Delta}$ converges weakly* in $\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$, and the limit can only be $x\mapsto\partial_{x}H(x,k)$. We can now conclude with $C$. Write $C=C_{1}+C_{2}$, where
$C_{1}$ $\coloneq-\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}\mathop{\rm sgn}(u_{j}^{n}-k)(F_{\mathrm{int}}^{j+1/2}(k,k)-F_{\mathrm{int}}^{j-1/2}(k,k))(\varphi_{j}^{n-1}-\varphi_{j}^{n})\Delta t$
$C_{2}$ $\coloneq-\sum_{n=1}^{+\infty}\sum_{j\in\mathbb{Z}}\int_{I_{j}}\mathop{\rm sgn}(u_{j}^{n}-k)F_{\Delta}^{\prime}(x)\varphi_{j}^{n}\Delta t.$
@@
${\left|C_{1}\right|}\leq 2TR\left(2\mu+\sup_{x\in\mathbb{R}}{\left|\partial_{x}H(x,k)\right|}\right){\left\|\partial_{t}\varphi\right\|}_{\mathbf{L}^{\infty}(\mathbb{R}^{+}\times\mathbb{R})}\Delta t.$
@@
for any k[$\k\in[\underline{u},\overline{u}]\backslash\mathcal{L}$. Then, repeat the argument from [KRT2004] to extend it for all $k\in[\underline{u},\overline{u}]$.